OpenStudy (anonymous):

Can someone help me determine if an improper integral converges or diverges?

OpenStudy (anonymous):

Hey

OpenStudy (anonymous):

What's your question?

OpenStudy (anonymous):

ok I will provide the example :D just give me a sec

OpenStudy (anonymous):

\[\int\limits_{1}^{\infty}(x ^{2}-6x+1)/(x ^{2}+4)\]

OpenStudy (anonymous):

I seriously dont understand the whole concept :P

OpenStudy (anonymous):

No, it doesn't.

OpenStudy (anonymous):

can you explain ur reasoning?

OpenStudy (anonymous):

Sure. Let's try the integral \[\int_{1}^\infty \frac{1}{x^2} dx\] We define improper integrals as \[ \lim_{a \rightarrow \infty} \int_{1}^a \frac{1}{x^2} dx \] Can you find this integral? Don't worry about the limit yet.

OpenStudy (anonymous):

= -x^-1

OpenStudy (anonymous):

Evaluated at the endpoints gives us?

OpenStudy (anonymous):

-1/a +1

OpenStudy (anonymous):

Perfect. Now apply the limit. Does it converge, and if so, to what?

OpenStudy (anonymous):

it converges it equals 1

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Jemurray?

OpenStudy (anonymous):

Sorry for the delay. Yes, that's correct. Now do the same for the integral \[\int_{1}^\infty \frac{1}{x} dx \]

OpenStudy (anonymous):

Well = lnIxI

OpenStudy (anonymous):

Okay, keep going :)

OpenStudy (anonymous):

lnIaI-lnI1I =lnIaI

OpenStudy (anonymous):

And if you apply the limit then we see it diverges

OpenStudy (anonymous):

Indeed it does. You seem to have it down pretty well. :)

OpenStudy (anonymous):

well this is easy but I am getting confused with this one

OpenStudy (anonymous):

I also dont know when a function is slowly hitting zero as it approaches infinity

OpenStudy (anonymous):

LIke how wld you know if it is approaching slowly or quickly?

OpenStudy (anonymous):

You just figured that out for yourself, though you may not realize it. We can find the "boundary" so to speak, or exactly how "fast" a function must approach zero such that its integral converges. Take \[ \int_{1}^\infty\frac{1}{ x^n} dx \] What must n be so that the integral will converge?

OpenStudy (anonymous):

N must be greater than 1. But i know that because my book tells me so

OpenStudy (anonymous):

what wld happen if n would be equal to 1?

OpenStudy (anonymous):

That was the second integral you calculated :)

OpenStudy (anonymous):

oh lol I see

OpenStudy (anonymous):

You see, if n is greater than one, then your boundary term will be 1/x^{something}, which approaches zero as x approaches infinity. If n is equal to one, we have a logarithm which diverges, and if n is less than one, the antiderivative has some power of x in the numerator, which obviously diverges.

OpenStudy (anonymous):

oh i see yes thats clear

OpenStudy (anonymous):

but lets return to the original problem

OpenStudy (anonymous):

How would i know like which part of the function dominates?

OpenStudy (anonymous):

What is the limit of the integrand as x approaches infinity?

OpenStudy (anonymous):

of which function?

OpenStudy (anonymous):

Your original problem.

OpenStudy (anonymous):

well there are so many x's in this problem. Some will be negative and others will be positive

OpenStudy (anonymous):

Try dividing top and bottom by x^2 and then taking the limit, ignoring terms that go to zero.

OpenStudy (anonymous):

sorry what do u mean?

OpenStudy (anonymous):

Well now i realized that when x approaches infinity x^2 dominates the numerator

OpenStudy (anonymous):

Dividing yields \[ \frac{1 - 6/x + 1/x^2}{1 + 4/x^2} \]

OpenStudy (anonymous):

the others are insignificant in comparison

OpenStudy (anonymous):

Yes, that's correct. So what's the limit?

OpenStudy (anonymous):

x^2/x^2=1 so it equals =x and so it diverges as x approaches infinty

OpenStudy (anonymous):

what i meant to say was that the integral of 1=x

OpenStudy (anonymous):

Well, I suppose you can think of it that way. It's not necessarily okay to just integrate the limit of the integrand but you can see that as x gets really really large you're effectively finding the area of the graph of y = 1, which is clearly infinite.

OpenStudy (anonymous):

So how wld you solve it?

OpenStudy (anonymous):

I wanna learn because i dont feel like i am doing it correctly. I am missing the basics in this area

OpenStudy (anonymous):

It's not a question of solving it or not. It doesn't converge, that's the end of the story.

OpenStudy (anonymous):

LOL hehe ok got it

OpenStudy (anonymous):

so for example \[\int\limits_{50}^{\infty}dx/x ^{3}\]

OpenStudy (anonymous):

This we can see automatically converges

OpenStudy (anonymous):

since as x becomes larger than 1/x^{3} is getting closer to 0

OpenStudy (anonymous):

No, because that would imply that 1/x converges too. Just remember what we said, it will converge if n > 1.

OpenStudy (anonymous):

well ok so if i had to explain a reasoning i wld say that /[1/x^{3}/] approaches 0 quickly as x approaches infinity?

OpenStudy (anonymous):

Quickly is a bad term. Instead, say that the antiderivative approaches zero, not the integrand itself.

OpenStudy (anonymous):

ohhhhh i seee fine

OpenStudy (anonymous):

Thanks for your help

OpenStudy (anonymous):

I am just trying to logically see this in my brain

OpenStudy (anonymous):

So that I will be able to solve on my own

OpenStudy (anonymous):

ok gn. If I have any more questions I will be back :D

OpenStudy (anonymous):

I hope u dont mind

OpenStudy (anonymous):

Sure, no problem , though I may be away tonight... if I can answer I will, if not, you can ask everyone else, or I'll be back tomorrow. Good luck :)

OpenStudy (anonymous):

ok Great thanks

OpenStudy (anonymous):

I finished my hmwrk So i am heading to bed :D

OpenStudy (anonymous):

I dont need anymore help :D